OpenStudy (anonymous):

can some help me with my homework on x and y intercepts.... PLEASE PLEASE PLEASE

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

its like 16 questions but they should be easy for you but i can do less if you have a limited time

OpenStudy (anonymous):

Maybe i can explain it so you'll know how to do a lot of them =D

OpenStudy (anonymous):

i know how to do them i just need like someone to check if i did them right

OpenStudy (anonymous):

Goody. Post 'em up

OpenStudy (anonymous):

1. x+y=4 2. -x+y=4 3. 2x+5y=10 do these please

OpenStudy (anonymous):

there is more by the way

OpenStudy (anonymous):

What are your answers?

OpenStudy (anonymous):

can you help me solve them

OpenStudy (anonymous):

ok for #1 i got 4 for both intercepts

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

That's right

OpenStudy (anonymous):

ok i dont know the rest can you help